Particulars

Brakepad

Gear

Direct Material

RM 8

RM 6

Direct Labour

24 hours @ RM0.25 per hour

16 hours @ RM 0.25 per hour

Selling Price

RM 25

RM 20

Variable Overhead

150% of direct labour

Fixed Overhead

RM 750

The directors want to be acquainted with the desirability of adopting any one of the following alternative sales mixes in the budget for the next period:

  1. 150 units of Brakepad and 350 units of Gear        
  2. 350 units of Brakepad and 150 units of Gear
  3. 400 units of Brakepad and 100 units of Gear
  4. 100 units of Brakepad and 400 units of Gear
  5. 250 units of Brakepad and 250 units of Gear
  6. 300 units of Brakepad and 300 units of Gear
  7. 500 units of Brakepad only
  8. 500 units of Gear only

REQUIRED;

  1. Compute the net profit for each of the above alternatives.

                                                                                                                                                 (24 marks)

  1. Choose the best alternative for the management.                                                               

                                                                                                                                                      (1 mark)
Answer #1

A Cost Volume Profit Analysis helps to determine the profitability of a product on the basis of change in volume of sales and costs. For this, it considers Sales, variable expenses, fixed expenses and the contribution margin.
Contribution margin is the difference between Sales and variable expenses. In this case variable expenses are direct materials, direct labor and variable overhead.

Direct materials per unit:
Brakepad = RM 8
Gear = RM 6

Direct labor per unit:
Brakepad = 24 hours @ RM0.25 per hour = 24 × 0.25 = RM 6

Gear = 16 hours @ RM0.25 per hour = 16 × 0.25 = RM 4

Variable overhead per unit:
Brakepad = 150% of direct labor = RM 6 × 150% = RM 9

Gear = 150% of direct labor = RM 4 × 150% = RM 6

a. A Contribution margin income statement is prepared to ascertain the net profit of each alternatives.

a. 150 units of Brakepad and 350 units of Gear CVP Analysis Brakepad Gear Total Units sold 150 units 350 units Selling price

d. 100 units of Brakepad and 400 units of Gear CVP Analysis Brakepad Gear Total Units sold 100 units 400 units Selling price

Total Gear O units 12,500 g. 500 units of Brakepad only CVP Analysis Brakepad Units sold 500 units Selling price per unit RM


a. The sales mix of 500 units of Gear only is the best alternative for the management as it produces a contribution margin of RM 2,000 and a net profit of RM 1,250 which is higher than all other alternatives.